Rewrite the equation -8x - 5y = -45
I know first I am to solve for y
5y + 5y cancels out the 5 and leaves y
-45 minus 5 = -50
So now my equation looks like this:
-8x = -50
Now I must solve for x
-8x divided in to -50 =
16
So would my answer be:
f(16) = -8 - 5
I am really stuck on these type of problems. Thank you for the wonderful help you give!!

Answer by AnlytcPhil(1806)

: Rewrite the equation -8x - 5y = -45
I know first I am to solve for y
5y + 5y cancels out the 5 and leaves y
-45 minus 5 = -50
So now my equation looks like this:
-8x = -50
Now I must solve for x
-8x divided in to -50 =
16
So would my answer be:
f(16) = -8 - 5
I am really stuck on these type of problems. Thank you for 
the wonderful help you give!!

No, I'm afraid this is all wrong.  Let's go thru it step by step:

-8x - 5y = -45

Get rid of any term on the left which does not contain a y. So we 
have to get rid of the first term, -8x.  To do this we add the 
opposite of -8x, namely +8x, to both sides

-8x - 5y = -45
+8x            + 8x
-------------------
     -5y = -45 + 8x

Now we divide every term by the coefficient of y, which is -5:

     -5y    -45     8x
    ————— = ———— + ————
      -5     -5     -5

In the first term on the left, the -5's cancel and leave just y

            -45     8x
        y = ———— + ————
             -5     -5

The first fraction on the right is just +9 since
-45 divided by -5 = +9

                 8x
        y = 9 + ————
                 -5

The second fraction with the negative sign on the bottom can be 
written with a minus sign in front

                 8x
        y = 9 - ————
                  5

You can leave it just like this, but the usual way is to write 
the second term as a fraction times x or

                8
        y = 9 - — x
                5
 
Also what is common is to write the term in x first.  So you 
might write that as


              8
        y = - — x + 9
              5
or
        y = -(8/5)x + 9

Edwin
AnlytcPhil@aol.com